# A Simple, Optimal and Efficient Algorithm for Online Exp-Concave Optimization

### v1

@@ 4.1 Restatement of the Open Problem
Open problem (Koren, 2013).
Under Assumption˜4 and $\mathcal{X}=\mathcal{B}(1)$, (a) Is it possible to find an SXO algorithm that attains the sample complexity of $\widetilde{O}(d/\epsilon)$ with only linear-in-$d$ runtime per iteration, i.e., $\widetilde{O}(d^{2}/\epsilon)$ runtime overall? (b) Is it possible to perform any better than $\widetilde{O}(d^{4}/\epsilon)$ runtime overall?
The first part of the open problem remains open. In particular, Mahdavi et al. (2015) shows an information-theoretic sample complexity lower bound of $\Omega(d/\epsilon)$ for SXO, which implies a runtime lower bound of $\Omega(d^{2}/\epsilon)$ since each gradient query costs $\Omega(d)$ time.
Section˜4.2 answers the second part in the affirmative by combining LightONS with the online-to-batch conversion, obtaining total runtime $\widetilde{O}(d^{3}/\epsilon)$ while maintaining the optimal $\widetilde{O}(d/\epsilon)$ sample complexity. In Section˜4.3, we present evidence that this $\widetilde{O}(d^{3}/\epsilon)$ runtime is likely unimprovable. If confirmed, it would refute the first part of the open problem and fully settle it.
4.2 Answering the Open Problem with LightONS
By applying the online-to-batch conversion for exp-concave functions from (Mehta, 2017), LightONS provides both high-probability and in-expectation convergence rate guarantees for SXO, as stated in the following theorem. The proof of the this theorem is provided in Section˜C.1.
Theorem 4.
Under Assumption˜4 and $\mathcal{X}=\mathcal{B}(1)$, let LightONS (Algorithm˜3) run for $T$ rounds with $T$ queries to the stochastic gradient oracle $\left\{\nabla f(\mathbf{x}_{t};\xi_{t})\right\}_{t=1}^{T}$, where $\left\{\mathbf{x}_{t}\right\}_{t=1}^{T}$ are the online decisions of LightONS, and let the aggregated solution be $\bar{\mathbf{x}}_{T}=\frac{1}{T}\sum_{t=1}^{T}\mathbf{x}_{t}$, for any $\delta\in(0,1)$, then the aggregated solution satisfies that, with probability at least $1-\delta$, with $T=\Theta\left(\frac{d}{\epsilon}\log\frac{d}{\epsilon}\log\frac{1}{\delta}\right)$,
$F(\bar{\mathbf{x}}_{T})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\leq\frac{1}{T}\left(\textsc{Reg}_{T}+4\sqrt{\frac{\textsc{Reg}_{T}}{2\gamma_{0}}\left(\log\frac{4\log T}{\delta}\right)}+\frac{8}{\gamma_{0}}\log\frac{4\log T}{\delta}\right)=O(\epsilon),$ (19)
where $\textsc{Reg}_{T}\leq\frac{d}{2\gamma_{0}}\log\left(1+\frac{G^{2}}{d\epsilon}T\right)+\frac{\gamma_{0}\epsilon D^{2}}{8}$ as in Theorem˜2. The total runtime of the algorithm is $O\left(\frac{d^{3}}{\epsilon}\log\frac{d}{\epsilon}\log\frac{1}{\delta}+\frac{d^{3.5}}{\sqrt{\epsilon}}\log\frac{d}{\epsilon}\log\frac{1}{\delta}\right)=\widetilde{O}\left(\frac{d^{3}}{\epsilon}\right)$. With $\frac{1}{\delta}=T^{\prime}=\Theta\left(\frac{d}{\epsilon}\log\frac{d}{\epsilon}\right)$, in expectation,
$\mathop{\mathbb{E}}\left[F(\bar{\mathbf{x}}_{T^{\prime}})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\right]\leq O(\epsilon).$ (20)
The total runtime of the algorithm is $O\left(\frac{d^{3}}{\epsilon}\log\frac{d}{\epsilon}+\frac{d^{3.5}}{\sqrt{\epsilon}}\log\frac{d}{\epsilon}\right)=\widetilde{O}\left(\frac{d^{3}}{\epsilon}\right)$.
4.3 Discussions on SXO
The LightONS-based SXO method of Theorem˜4 matches the best known SXO runtime up to polylogarithmic factors. An online-to-batch conversion of OQNS yields the same asymptotical runtime $\widetilde{O}(d^{3}/\epsilon)$, but, as discussed in Section˜3, its regret bound carries large constants, which slows convergence in stochastic settings. Our experiments in Section˜3.4, conducted in SXO settings, confirm this difference.
We conjecture that no SXO algorithm can asymptotically beat total runtime $\widetilde{O}(d^{3}/\epsilon)$. Two observations support this view:
  • Fast matrix-multiplication. Linear regression with random design, a special case of SXO, reduces to solving a linear system with $d$ variables and $\widetilde{O}(d/\epsilon)$ equations. Although fast matrix-multiplication accelerates this to $\widetilde{O}(d^{\omega}/\epsilon)$ time (Ibarra et al., 1982), such techniques do not plausibly extend beyond linear models. In practice, the runtime effectively reverts to $\widetilde{O}(d^{3}/\epsilon)$.
  • Offline convex optimization. SXO methods based on Empirical Risk Minimization (ERM) (Koren and Levy, 2015; Mehta, 2017) likewise fail to surpass the $\widetilde{O}(d^{3}/\epsilon)$ barrier even with cutting-edge offline solvers (Lee et al., 2015; Jiang et al., 2020), as detailed in Section˜C.2. Moreover, such methods require $\Omega(d^{2}+d/\epsilon)$ working memory to store all samples and well-roundedness of the domain, whereas the OXO-based algorithms typically require only $O(d^{2})$ memory and no such geometric assumptions.
5 Applications to Various Problems
ONS radiates influence well beyond the classical settings of OXO and SXO. By preserving the elegant algorithmic structure of ONS, LightONS can be seamlessly integrated into various applications where ONS serves as a computational core to substantially improve efficiency. In this section, we elaborate on three representative applications of ONS where LightONS fit in without compromising statistical advantages or requiring much new analytical efforts.
  1. (i)
    Gradient-norm adaptivity. ONS achieves a regret bound scaling with the accumulated squared gradient norms, which further implies small-loss adaptivity in OXO and comparator-norm adaptivity in OCO.
  2. (ii)
    Parametric stochastic bandits. Instead of directly minimizing the regret, ONS serves as the parameter estimator to handle customized local norms that is crucial for the exploration-exploitation trade-off.
  3. (iii)
    Memory-efficient OXO. When the online gradients exhibit low-rank structure, combining ONS with matrix sketching can achieve linear-in-$d$ runtime and working memory.
We note that in these scenarios, OQNS can hardly be applied due to its significant algorithmic deviations from disciplined online learning frameworks, especially its log-barrier regularization.
5.1 Gradient-Norm Adaptivity
$\textsc{Reg}_{T}(\mathbf{u})=O\left(d\log G_{T}\right),\quad\text{where}~G_{T}\triangleq\sum_{t=1}^{T}\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2}.$ (21)
In benign environments, $G_{T}$ can be $O(1)$, for example, when the loss is nearly fixed, decisions converge quickly, yielding regret far below worst-case bounds that scale with $T$. Yet, since $G_{T}\leq G^{2}T$, the gradient-norm adaptive bound safeguards the minimax-optimality in $T$. Prior works have leveraged Eq.˜21 to achieve such stronger types of adaptivity, including small-loss adaptivity for OXO with smoothness and comparator-norm adaptivity for unbounded OCO, as discussed below.
OXO with smoothness.
Orabona et al. (2012) demonstrate that ONS directly implies the desired regret bound in Eq.˜21 for OXO. Under smoothness assumptions, the gradient-norm adaptivity can further be improved to small-loss adaptivity (Srebro et al., 2010; Zhao et al., 2020), where the regret scales with the cumulative loss of the best comparator in hindsight, regardless of online decisions. This is formalized as follows:
$\textsc{Reg}_{T}=O\left(d\log L_{T}\right),\quad\text{where}~L_{T}\triangleq\min_{\mathbf{u}\in\mathcal{X}}\sum_{t=1}^{T}\left(f_{t}(\mathbf{u})-\min_{\mathbf{x}\in\mathcal{X}}f_{t}(\mathbf{x})\right),$ (22)
when the loss function $f_{t}:\mathcal{X}\to\mathbb{R}$ is $H$-smooth, i.e., for any $t\in[T]$ and any pair $(\mathbf{x},\mathbf{y})\in\mathcal{X}^{2}$, $\left\|\nabla f(\mathbf{x})-\nabla f(\mathbf{y})\right\|_{2}\leq H\left\|\mathbf{x}-\mathbf{y}\right\|_{2}$.
Unbounded OCO.
Cutkosky and Orabona (2018) extend the gradient-norm adaptive OXO regret in Eq.˜21 to a comparator-norm adaptive OCO regret, i.e.,
$\textsc{Reg}_{T}(\mathbf{u})=\widetilde{O}\left(\left\|\mathbf{u}\right\|_{2}\sqrt{dG_{T}}\right),\quad\text{for any (unbounded) comparator}~\mathbf{u}\in\mathbb{R}^{d},$ (23)
without any prior knowledge of $\mathbf{u}$. Such guarantees for unbounded comparators are unachievable via classical OCO algorithms that explicitly depend on the domain diameter $D$, for example, OGD demands an explicit $D$ to obtain $O(DG\sqrt{T})$ regret (Zinkevich, 2003; Abernethy et al., 2008).
Nevertheless, both (Orabona et al., 2012; Cutkosky and Orabona, 2018) incur the worst-case $\widetilde{O}(d^{\omega}T)$ runtime bottleneck of ONS. Below, we show that LightONS can replace ONS in these algorithms to substantially improve efficiency without compromising statistical guarantees.
Improvements by LightONS.
LightONSrecovers ONS’s gradient-norm adaptive bound in Eq.˜21, thereby leading to small-loss adaptivity for OXO under smoothness assumptions, as well as comparator-norm adaptivity for unbounded OCO. We state the theoretical guarantees based on LightONS and omit those of the original ONS-based methods as they are identical except for the runtime. The proofs are deferred to Section˜D.1.
Theorem 5 (LightONS’s improvement for OXO with smoothness).
where $L_{T}$ is defined in Eq.˜22. The total runtime of the algorithm is $O((\texttt{EP}_{\mathcal{X}}+d^{2})T+d^{\omega}\sqrt{T}\log T)$.
Theorem 6 (LightONS’s improvement for unbounded OCO).
$\textsc{Reg}_{T}(\mathbf{u})\leq\widetilde{O}\left(\sqrt{d\sum_{t=1}^{T}\left(\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}\mathbf{u}\right)^{2}}\right)\leq\widetilde{O}\left(\left\|\mathbf{u}\right\|_{2}\sqrt{dG_{T}}\right).$
The total runtime of the algorithm is $O(d^{2}T+d^{\omega}\sqrt{T}\log T)$.
Specifically, Theorem˜5 recovers Theorem 1 of Orabona et al. (2012), and Theorem˜6 recovers Theorem 8 of Cutkosky and Orabona (2018). The computational gains from replacing ONS with LightONS, discussed earlier, apply here.
Comparison with OQNS.
We remark that OQNS can hardly achieve full gradient-norm adaptivity, as the log-barrier regularization introduces an unavoidable $O(\log T)$ bias term independent of the gradient norm. From the perspective of OMD, the regret can be decomposed into a stability term and a bias term. Adopted from (Hazan, 2016), the regret decomposition of a typical OMD algorithm is as follows:
While ONS and LightONS, as instances of OMD, can flexibly balance stability and bias, OQNS sacrifices such flexibility for computational efficiency. OQNS employs a highly stable log-barrier, which suppresses the stability term but inflates the bias in a manner resistant to small gradient norms. Specifically, the log-barrier contributes a term like $\log(D^{2}-\left\|\mathbf{u}\right\|_{2}^{2})$ in the regret bound for any comparator $\mathbf{u}\in\mathcal{B}(D)$. As $\left\|\mathbf{u}\right\|_{2}$ approaches $D$, this term diverges. This issue is mitigated using the fixed-share trick (Mhammedi and Gatmiry, 2023), an illustration of which is as follows:
$\sum_{t=1}^{T}\left(f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{u})\right)\leq\underbrace{\sum_{t=1}^{T}\left(f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{v})\right)}_{\text{fixed-share regret}}+\underbrace{\sum_{t=1}^{T}\left(f_{t}(\mathbf{v})-f_{t}(\mathbf{u})\right)}_{\text{fixed-share margin}},\quad\text{where}~\mathbf{v}=\left(1-\frac{1}{T}\right)\mathbf{u}.$
In the fixed-share margin term, Lipschitzness yields $f_{t}(\mathbf{v})-f_{t}(\mathbf{u})\leq O(GD/T)$, summing to $O(1)$ over $T$ rounds. In the fixed-share regret term, since $1-\left\|\mathbf{v}\right\|_{2}^{2}\geq\Omega(D^{2}/T)$, the term $\log(D^{2}-\left\|\mathbf{v}\right\|_{2}^{2})$ contributes $O(\log T)$ to the regret, which is independent of $G_{T}$ and prevents full gradient-norm adaptivity.
5.2 Logistic Bandits
Generalized linear bandits (GLB) constitutes an important class of parametric stochastic bandits, where the expected loss depends on an unknown parameter through a known link function. This formulation introduces non-linearities into linear bandits, enhancing their expressivity while retaining tractable structures. We focus on logistic bandits as a representative instance of GLB to concretely illustrate the applicability of ONS and the integration of LightONS.
Problem setting.
Logistic bandits can be interpreted as interactions between a learner and an adversary, which unfolds as follows: At each round $t\in[T]$, the learner selects an arm $\mathbf{x}_{t}$ from an arm set $\mathcal{X}_{t}\subseteq\mathcal{B}(1)^{K}$ with $K$ arms, and suffers a stochastic loss $y_{t}\in\left\{0,1\right\}$, where $\mathop{\mathbb{P}}\left(y_{t}=1\,|\,\mathbf{x}_{t}\right)=\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}^{\star})$. The true parameter $\mathbf{w}^{\star}\in\mathcal{W}=\mathcal{B}(D/2)$ is unknown, and $\sigma(z)=1/(1+\exp(-z))$ denotes the sigmoid function. The performance of the learner is measured by its pseudo-regret, quantifying the cumulative expected loss against the optimal arms in hindsight, which is defined as
A challenge in GLB lies in its dependence on $\kappa=\max_{\mathbf{x}\in\mathcal{X},\mathbf{w}\in\mathcal{W}}1/\sigma^{\prime}(\mathbf{x}^{\mathsf{T}}\mathbf{w})$, a problem-dependent constant that may grow expoenentially with $D$. Directly applying ONS to GLB yields a pseudo-regret bound of $O(\kappa\sqrt{T})$, which becomes vacuous for $\kappa=\Omega(\sqrt{T})$ (Jun et al., 2017).
Current results.
Recently, Zhang et al. (2025) propose the first GLB algorithm that achieves $\kappa$-free-leading-term pseudo-regret with constant working memory and constant per-round time. Their pseudo-regret upper bound is $O(d\sqrt{T}\log T+\kappa(d\log T)^{2})$, which remains sublinear in $T$ for $\kappa=o(T)$.
Their key idea is to estimate the parameter $\mathbf{w}^{\star}$ with a “look-ahead” variant of ONS, whose analytical properties can remove the dependence on $\kappa$ in the pseudo-regret’s leading terms while retaining the constant-memory and constant-time efficiency of ONS. An illustration of their “look-ahead” ONS is provided as follows, where $\ell_{t}(\mathbf{w})=-y_{t}\log\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w})-(1-y_{t})\log(1-\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}))$ denotes the logistic loss for parameter estimation at round $t$:
$\widetilde{H}_{t}=\epsilon I+\sum_{i=1}^{t-1}\nabla^{2}\ell_{i}(\mathbf{w}_{i+1})+\eta\nabla^{2}\ell_{t}(\mathbf{w}_{t}),~\widehat{\mathbf{w}}_{t+1}=\mathbf{w}_{t}-\frac{1}{\eta}\widetilde{H}_{t}^{-1}\nabla\ell_{t}(\mathbf{w}_{t}),~\mathbf{w}_{t+1}=\Pi^{\widetilde{H}_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{w}}_{t+1}].$ (25)
As a result, their algorithm incurs the worst-case $\widetilde{O}((d^{2}K+d^{\omega})T)$ runtime bottleneck of ONS, where $K$ is the number of arms. Below we show that plugging LightONS can improve the runtime while preserving the $\kappa$-free-leading-term pseudo-regret guarantee.
Improvements by LightONS.
The LightONS counterpart of the “look-ahead” ONS replaces the Mahalanobis projection in Eq.˜25 with the projection-hysteresis mechanism as follows:
$\mathbf{w}_{t+1}=\begin{cases}\widehat{\mathbf{w}}_{t+1}&\text{if}~\left\|\widehat{\mathbf{w}}_{t+1}\right\|_{2}\leq kD/2\\ \Pi^{\widetilde{H}_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{w}}_{t+1}]&\text{otherwise}\end{cases}.$ (26)
The replacement entails theoretical guarantees as follows, with proofs deferred to Section˜D.2.
Theorem 7 (LightONS’s improvement for (Zhang et al., 2025)).
The algorithm in Theorem˜7 reduces the total runtime from $\widetilde{O}((d^{2}K+d^{\omega})T)$ to $\widetilde{O}(d^{2}KT+d^{\omega}\cdot\min\{\sqrt{\kappa dT},T\})$, while preserving the $\kappa$-free-leading-term pseudo-regret guarantee. When $\kappa=o(T)$, as required for sublinear pseudo-regret, the improved algorithm is asymptotically faster than the original algorithm.
We remark that that Eq.˜26 does not suffer from improper learning concerns, because the algorithm’s decision is the arm $\mathbf{x}_{t}$, rather than the estimated parameter $\mathbf{w}_{t}$. Consequently, the replacement is free from the improper-to-proper conversion and resembles LightONS.Core.
Comparison with OQNS.
OQNS can hardly be integrated into $\kappa$-free-leading-term GLB algorithms, as it is tailored to the OXO protocol and lacks the flexibility to handle customized local norms beyond those of ONS. Specifically, the local norm $\widetilde{H}_{t}$ in (Zhang et al., 2025), i.e., Eq.˜25, is not a simple accumulation of gradient outer products, unlike the OQNS’s local norm in Eq.˜7b. In contrast, LightONS retains the structural flexibility of ONS, enabling integration into GLB with minimal modifications.
5.3 Memory-Efficient OXO
In addition to the projection issue, another challenge of ONS is its $O(d^{2})$ working memory, in contrast to OGD’s $O(d)$ working memory. Luo et al. (2016) propose Sketched Online Newton Step (SON), which mitigates this issue by incorporating matrix sketching into ONS.
Current results.
SON achieves runtime and working memory linear in $d$ when the sketched dimension $d^{\prime}\ll d$. The value of $d^{\prime}$ typically depends on the problem’s intrinsic dimensionality, such as the number of non-zero eigenvalues in the Hessian-related matrix. The theoretical guarantees of SON are summarized in the following proposition.
Proposition 3 (Theorem 3 of Luo et al. (2016)).
Under Assumption˜2 and additional assumptions described in Eqs.˜30 and 31, SON (Algorithms 1 and 6 of Luo et al. (2016)) satisfies that,
$\textsc{Reg}_{T}(\mathbf{u})\leq\frac{d^{\prime}}{\gamma}\log\left(1+\frac{G^{2}}{2d^{\prime}\epsilon}T\right)+\frac{\gamma\epsilon D^{2}}{8}+\frac{\Delta_{1:T}}{2\gamma},$ (27)
where $\mathbf{u}$ and $\gamma$ are restricted by Eq.˜31, $\epsilon$ is the preconditioner coefficient similar to that in ONS and LightONS, and $\Delta_{1:T}$ is the cumulative sketching error, which can be bounded as
$\Delta_{1:T}\leq\min_{j\in[d^{\prime}]}\frac{2d^{\prime}}{(d^{\prime}-j+1)\epsilon}\sum_{i=j}^{d}\lambda_{i}\left(\sum_{t=1}^{T}\nabla f_{t}(\mathbf{x}_{t})\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}\right),$ (28)
and $\lambda_{i}$ is the $i$-th greatest eigenvalue. The working memory is $O(d^{\prime}d)$ and the total runtime is  [^3] The $\log T$ term in Eq. 29 arises from the eigendecomposition underlying sketching. Although Theorem 3 of Luo et al. (2016) ignores it, we include it for comparison with LightONS and OQNS.
$O\left(d^{\prime}dT\log T\right).$ (29)
However, SON demands additional assumptions. Its domain $\mathcal{X}_{t}$ must be an intersection of two parallel half-spaces, onto which Mahalanobis projections admit closed-form with $O(d^{2})$ time. For bounded convex domains in OXO, SON loses its computational advantage and reverts to the high computational cost of ONS. The domain restriction further imposes a stronger assumption on loss functions. Under Assumptions˜1, 2 and 3, the curvature parameter $\gamma_{0}=\frac{1}{2}\min\{\frac{1}{DG},\alpha\}$ in Lemma˜1 collapses to zero for the unbounded intersection-of-parallel-half-spaces domain, reducing exp-concavity to convexity. Consequently, SON assumes an explicit quadratic property analogous to Lemma˜1, rather than the standard Assumption˜3. We list these additional assumptions as follows.
    $\mathcal{X}_{t}=\{\mathbf{x}\,|\,|\mathbf{w}_{t}^{\mathsf{T}}\mathbf{x}|<D/2\},\quad\text{where}~\mathbf{w}_{t}\in\mathbb{R}^{d}~\text{is known and}~\left\|\mathbf{w}_{t}\right\|=1.$ (30)
  • Additional assumption on loss functions. For any $t\in[T]$, the loss function $f_{t}:\mathcal{X}\to\mathbb{R}$, the trajectory $\{\mathbf{x}_{t}\}_{t=1}^{T}$ and the comparator $\mathbf{u}\in\bigcap_{t=1}^{T}\mathcal{X}_{t}$ satisfies that
Improvements by LightONS.
Replacing all accesses to the Hessian-related matrix $A_{t}$ with the sketching primitives as in SON yields a hybrid method that combines the projection efficiency of LightONS with the memory efficiency of SON. This hybrid method extends linear-in-$d$ runtime and working memory to the standard OXO setting in Section˜2.1, while retaining the regret $O(d^{\prime}\log T)$ of SON. Theoretical guarantees of the hybrid method are summarized in the following theorem, with algorithms and proofs deferred to Section˜D.3.
Theorem 8 (LightONS’s improvement for memory-efficient OXO).
Under Assumptions˜1, 2 and 3, there exists an algorithm, which satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})\leq\frac{d^{\prime}}{\gamma_{0}}\log\left(1+\frac{G^{2}}{2d^{\prime}\epsilon}T\right)+\frac{\gamma_{0}\epsilon D^{2}}{8}+\frac{\Delta_{1:T}}{2\gamma_{0}},$ (32)
where $\gamma_{0}=\frac{1}{2}\min\left\{\frac{1}{DG},\alpha\right\}$, $\epsilon$ is the preconditioner coefficient similar to that in LightONS, $\Delta_{1:T}$ is the cumulative sketching error, which can be bounded as
$\Delta_{1:T}\leq\min_{j\in[d^{\prime}]}\frac{2d^{\prime}}{(d^{\prime}-j+1)\epsilon}\sum_{i=j}^{d}\lambda_{i}\left(\sum_{t=1}^{T}\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}\right),$ (33)
$\nabla g_{t}(\mathbf{y}_{t})$ is the surrogate gradient as in Lemma˜3, and $\lambda_{i}$ is the $i$-th greatest eigenvalue. The working memory is $O(d^{\prime}d)$ and the total runtime is
$O\left(\left(\texttt{EP}_{\mathcal{X}}+d^{\prime}d\log T\right)T+d^{\omega}\sqrt{(d+\Delta_{1:T})T/\epsilon}\log T\right).$ (34)
Comparison with OQNS.
Mhammedi and Gatmiry (2023) mention the possibility of combining OQNS with sketching but provide neither algorithms nor analysis. Incorporating sketching further complicates OQNS’s already intricate analysis, as the sketching error interacts with the log-barrier, Hessian approximation, and decisions’ update, as shown in Eqs.˜7b and 7a. In contrast, within the OMD framework, sketching errors naturally appear as an additive term in the regret bound via OMD’s stability term, illustrated in Eq.˜24, since OMD’s regret decomposition is somewhat orthogonal to sketching.
6 Conclusion
For online exp-concave optimization (OXO), we propose LightONS, a simple yet powerful variant of ONS. By combining the projection-hysteresis mechanism with the improper-to-proper conversion, LightONS achieves significant reductions in worst-case total runtime while preserving the regret optimality of ONS. These gains extend to the stochastic optimization setting, answering a COLT’13 open problem on efficient and optimal SXO (Koren, 2013). Moreover, due to its fidelity to the mirror-descent udpate of ONS, LightONS serves as a drop-in replacement across diverse applications, including regret bounds adaptive to gradient norms, parametric stochastic bandits, and memory-efficient settings, all without compromising statistical guarantees.
Several important directions remain open. First, while LightONS adapts to gradient norms, achieving adaptivity to gradient variation for OXO within $\widetilde{O}(d^{2}T)$ time remains an open challenge (Chiang et al., 2012; Zhao et al., 2024). Second, applying the LightONS technique to other preconditioned online learning algorithms may not improve their (asymptotic) runtime if the bottleneck lies in other operations, such as explicit matrix factorization rather than Mahalanobis projections (e.g., in AdaGrad (Duchi et al., 2011)). Accelerating these algorithms remains unclear. These future directions motivate the development of broader, general-purpose acceleration techniques for Hessian-related online learning algorithms.
Appendix A Technical Lemmas
In this section, we present technical lemmas for the analysis of the proposed OXO algorithms.
A.1 Elliptical Potential Lemmas
We begin with two lemmas addressing Hessian-related matrices, which are essential for both regret and runtime analysis. Lemma˜5 is standard in the ONS literature (e.g., [Hazan et al., 2007, Luo et al., 2016]). The proof of Lemma˜6 mirrors that of Lemma˜5, differing only in the potential function: $F(A)=-\operatorname{tr}(A)$ versus $F(A)=\log\operatorname{det}(A)$.
Lemma 5 (elliptical potential lemma).
@@ A.1 Elliptical Potential Lemmas
Proof of Lemmas˜5 and 6.
The first equality follows from $A_{i}-A_{i-1}=\mathbf{v}_{i}\mathbf{v}_{i}^{\mathsf{T}}$ (by the definition of $A_{i}$) and $\mathbf{u}^{\mathsf{T}}A\mathbf{v}=\left\langle A,\mathbf{v}\mathbf{u}^{\mathsf{T}}\right\rangle_{\operatorname{F}}$ (by the definition of the $\left\langle\cdot,\cdot\right\rangle_{\operatorname{F}}$); The inequality from the concavity of $F$; The final equality from telescoping. Substituting the respective potentials into Eq.˜37 yields Lemmas˜5 and 6. Concretely,
  • For Lemma˜5, $F(A)=\log\operatorname{det}(A)$ and $\nabla{F}(A)=A^{-1}$.
  • For Lemma˜6, $F(A)=-\operatorname{tr}(A^{-1})$ and $\nabla{F}(A)=A^{-2}$.
A.2 Proof of Lemma˜1

### v2

@@ B.3 Numerical Lemmas
s.t. $\mathbf{x}^{\mathsf{T}}\mathbf{x}\leq R^{2}$
The Lagrangian of this quadratic program is $\mathcal{L}=(\mathbf{x}-\mathbf{y})^{\mathsf{T}}A(\mathbf{x}-\mathbf{y})+\mu(\mathbf{x}^{\mathsf{T}}\mathbf{x}-R^{2})$. According to the KKT conditions, $\nabla\mathcal{L}=2A(\mathbf{x}-\mathbf{y})+2\mu\mathbf{x}=\bm{0}$ with $\mu>0$. (Otherwise, $\mu=0$ implies $\mathbf{y}\in\mathcal{B}(R)$ and the projection is trivial.) Rearranging $\nabla\mathcal{L}=\bm{0}$ yields $\mathbf{x}=(A+\mu I)^{-1}A\mathbf{y}$, thus $\mathbf{x}^{\mathsf{T}}\mathbf{x}=R^{2}$ is equivalent to $\rho(\mu)=0$, which proves Eq.˜17.
To prove Eq.˜18, since $\rho$ monotonically decreases on $0,\infty)$, it suffices to bound $\rho(\mu)$. By the orthogonality of $Q$, we have $\sum_{i=1}^{d}v_{i}^{2}=\left\|Q\mathbf{y}\right\|_{2}^{2}=\left\|\mathbf{y}\right\|_{2}^{2}$. Therefore,
Then it is straightforward that $\rho((\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{d})\geq 0$ and $\rho((\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{1})\leq 0$, yielding that $\mu^{\ast}\in((\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{d},(\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{1})$ and $\mu^{\ast}$ is unique on $0,\infty)$. ∎
@@
Let $\mathbf{x}^{\ast}=(A+\mu^{\ast}I)^{-1}A\mathbf{y}$, where $\rho(\mu^{\ast})=0$ as in Lemma˜8. If $\widetilde{\mu}\geq 0$ and $\left|\widetilde{\mu}-\mu^{\ast}\right|<\frac{\lambda_{d}}{\left\|\mathbf{y}\right\|_{2}}\zeta$, then $\mathbf{x}\in\mathcal{B}(R)$ and $\left\|\mathbf{x}-\mathbf{x}^{\ast}\right\|_{2}\leq\zeta$, where $\mathbf{x}$ is constructed from $\widetilde{\mu}$ as
Proof of Lemma˜9.
Eq.˜19 immediately implies $\mathbf{x}\in\mathcal{B}(R)$.
To prove $\left\|\mathbf{x}-\mathbf{x}^{\ast}\right\|_{2}\leq\zeta$, since $\left\|\mathbf{x}-\mathbf{x}^{\ast}\right\|_{2}\leq\left\|\widetilde{\mathbf{x}}-\mathbf{x}^{\ast}\right\|_{2}$ by Lemma˜10, it suffices to prove $\left\|\widetilde{\mathbf{x}}-\mathbf{x}^{\ast}\right\|_{2}\leq\zeta$. By the definition of $v_{i}$ in Lemma˜8, we have
$\leq\sum_{i=1}^{d}v_{i}^{2}\left(1-\frac{1}{1+\frac{\left|\widetilde{\mu}-\mu^{\ast}\right|}{\lambda_{i}}}\right)^{2}\leq\sum_{i=1}^{d}v_{i}^{2}\left(\frac{\widetilde{\mu}-\mu^{\ast}}{\lambda_{i}}\right)^{2}\leq\left\|\mathbf{y}\right\|_{2}^{2}\left(\frac{\widetilde{\mu}-\mu^{\ast}}{\lambda_{d}}\right)^{2}<\zeta^{2}.$
The first inequality uses that $\lambda_{i}>0$, $\widetilde{\mu}\geq 0$ and $\mu^{\ast}\geq 0$; The second inequality comes from the following inequality, $(1-\frac{1}{1+a})^{2}=\frac{a^{2}}{(a+1)^{2}}\leq a^{2}$ for any $a\geq 0$; The third inequality uses that $\sum_{i=1}^{d}v_{i}^{2}=\left\|\mathbf{y}\right\|_{2}^{2}$; The last inequality uses that $|\widetilde{\mu}-\mu^{\ast}|<\frac{\lambda_{d}}{\left\|\mathbf{y}\right\|_{2}}\zeta$. ∎
With the help of these lemmas, we can prove Lemma˜7.
Proof of Lemma˜7.
First, we justify the bisection. By Lemma˜8, the wanted zero satisfies that $\mu^{\ast}\in(a_{1},b_{1})=((\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{d},(\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{1})$, which implies the initial interval of bisection (Line 3 in Algorithm˜4). Furthermore, $\rho$ monotonically decreases on $0,\infty)$, which implies the selection of $a_{t+1}$ and $b_{t+1}$ (Lines 4–6 in Algorithm˜4).
Then by Lemma˜9, Algorithm˜4 achieves an error $\left\|\mathbf{x}-\mathbf{x}^{\ast}\right\|_{2}\leq\zeta$.
Finally, we bound the runtime. We note that choice 1 and 2 are equivalent, since
With choice 1, each iteration requires $d^{\omega}$ arithmetic operations, where the computational bottleneck lies in the matrix inversion $(A+\mu I)^{-1}$, resulting in a runtime of $O(d^{\omega}n)$. With choice 2, the tridiagonalization of $A$ requires $O(d^{3})$ arithmetic operations [Golub and Van Loan, 2013], and each iteration can evaluate $(C+\mu I)^{-1}\mathbf{w}$ via the Thomas algorithm with only $O(d)$ arithmetic operations [Golub and Van Loan, 2013], resulting in a runtime of $O(d^{3}+dn)$. ∎
Appendix C Proofs for Section˜3
In this section, we prove the theoretical guarantees of LightONS.
C.1 Proof of Lemma˜2